# EraseDiff: Erasing Data Influence in Diffusion Models

### v3

@@ 3.2 Solution
The objective aims to minimize $\mathcal{L}_{r}({\bm{\theta}};{\mathcal{D}}_{r})+\lambda g({\bm{\theta}};\bm{\phi}^{\ast}({\bm{\theta}}))$, let ${\bm{\theta}}^{\ast}$ be the optimal solution to this objective. Note that $g({\bm{\theta}},\bm{\phi}^{\ast}({\bm{\theta}}))={\mathcal{L}}_{f}({\bm{\theta}};{\mathcal{D}}_{f})-\operatorname{min}{\mathcal{L}}_{f}(\bm{\phi}^{\ast}({\bm{\theta}});{\mathcal{D}}_{f})\geq 0$ as $\bm{\phi}$ starts from ${\bm{\theta}}$ and is update to decreas ${\mathcal{L}}_{m}(\bm{\phi};{\mathcal{D}}_{f})$. This will decrease to 00 for minimizing the above objective. Therefore, at the optimal solution ${\bm{\theta}}^{\ast}$, we have $g({\bm{\theta}}^{\ast},\bm{\phi}^{\ast}({\bm{\theta}}^{\ast}))=0$. This further implies that ${\mathcal{L}}_{f}({\bm{\theta}}^{*};{\mathcal{D}}_{f})=\operatorname{min}{\mathcal{L}}_{f}(\bm{\phi}^{\ast}({\bm{\theta}}^{*});{\mathcal{D}}_{f})$, meaning that ${\bm{\theta}}^{*}$ is the current optimal solution of ${\mathcal{L}}_{f}({\bm{\theta}};D_{f})$ because we cannot update further the optimal solution. Moreover, we have ${\bm{\theta}}^{*}$ as the local minima of $\mathcal{L}_{r}({\bm{\theta}};{\mathcal{D}}_{r})$ in sufficiently small vicinity considered, because in the small vicinity around ${\bm{\theta}}^{*}$, $g\left({\bm{\theta}},\bm{\phi}^{*}({\bm{\theta}}^{*})\right)=0$ provides no further improvements for the above sum, any increase in the above objective in the vicinity of ${\bm{\theta}}^{*}$ would primarily be due to an increase in $\mathcal{L}_{r}({\bm{\theta}};{\mathcal{D}}_{r})$.
4 Related Work
Memorization. Privacy of generative models has been studied extensively for GANs (Feng et al., 2021; Meehan et al., 2020; Webster et al., 2021) and generative language models (Carlini et al., 2022, 2021; Jagielski et al., 2022; Tirumala et al., 2022). These generative models often risk replicating from their training data. Recently, several studies (Carlini et al., 2023; Somepalli et al., 2023b, a; Vyas et al., 2023) investigated these data replication behaviors in diffusion models, raising concerns about the privacy and copyright issues. Possible mitigation strategies are deduplicating and randomizing conditional information (Somepalli et al., 2023b, a), or training models with differential privacy (DP) (Abadi et al., 2016; Dwork et al., 2006; Dwork, 2008; Dockhorn et al., 2022). However, leveraging DP-SGD (Abadi et al., 2016) may cause training to diverge (Carlini et al., 2023).
Malicious misuse. Diffusion models usually use training data from varied open sources and when such unfiltered data is employed, there is a risk of it being tainted (Chen et al., 2023b) or manipulated (Rando et al., 2022), resulting in inappropriate generation (Schramowski et al., 2023). They also risk the imitation of copyrighted content, e.g., mimicking the artistic style (Gandikota et al., 2023a; Shan et al., 2023). To counter inappropriate generation, data censoring (Gandhi et al., 2020; Birhane & Prabhu, 2021; Nichol et al., 2021; Schramowski et al., 2022) where excluding black-listed images before training, and safety guidance where diffusion models will be updated away from the inappropriate/undesired concept (Gandikota et al., 2023a; Schramowski et al., 2023) are proposed. Shan et al. (2023) propose protecting artistic style by adding barely perceptible perturbations to the artworks before public release. Yet, Rando et al. (2022) argue that DMs can still generate content that bypasses the filter. Chen et al. (2023b) highlight the susceptibility of DMs to poison attacks, where target images are generated with specific triggers.
Machine unlearning. Removing data directly involves retraining the model from scratch, which is inefficient and impractical. Thus, to reduce the computational overhead, efficient machines unlearning methods (Romero et al., 2007; Karasuyama & Takeuchi, 2010; Cao & Yang, 2015; Ginart et al., 2019; Bourtoule et al., 2021; Wu et al., 2020; Guo et al., 2020; Golatkar et al., 2020; Mehta et al., 2022; Sekhari et al., 2021; Chen et al., 2023a; Tarun et al., 2023b) have been proposed. Several studies (Gandikota et al., 2023a, b; Heng & Soh, 2023a, b; Fan et al., 2023; Zhang et al., 2023; Bui et al., 2024) recently introduce unlearning in diffusion models. Most of them (Gandikota et al., 2023a, b; Heng & Soh, 2023a; Zhang et al., 2023) mainly focus on text-to-image models and high-level visual concept erasure. Heng & Soh (2023b) adopt Elastic Weight Consolidation (EWC) and Generative Replay (GR) from continual learning to perform unlearning effectively without access to the training data. Heng and Soh’s method can be applied to a wide range of generative models, however, it needs the computation of FIM for different datasets and models, which may lead to significant computational demands. Fan et al. (2023) propose a very potent unlearning algorithm called SalUn that shifts attention to important parameters w.r.t. the forgetting data. SalUn can perform effectively across image classification and generation tasks.
In this work, we introduce a simple yet effective unlearning algorithm for diffusion models by formulating the problem as a constrained optimization problem. Below, we will show that our algorithm is not only faster than Heng and Soh’s method (Heng & Soh, 2023b) and Fan’s method (Fan et al., 2023), but even outperforms these methods in terms of the trade-off between the forgetting and preserving model utility.
5 Experiment
We evaluate EraseDiff in various scenarios, including removing images with specific classes/concepts, to answer the following research questions (RQs): (i) Can typical machine unlearning methods be applied to diffusion models? (ii) Is the proposed method able to remove the influence of ${\mathcal{D}}_{f}$ in the diffusion models? (iii) Is the proposed method able to preserve the model utility while removing ${\mathcal{D}}_{f}$? (iv) Is the proposed method efficient in removing the data? (v) How does the proposed method perform on the public well-trained models?
5.1 Setup
Experiments are reported on CIFAR10 (Krizhevsky et al., 2009) with DDPM, Imagenette (Howard & Gugger, 2020) with Stable Diffusion (SD) for class-wise forgetting, I2P (Schramowski et al., 2023) dataset with SD for concept-wise forgetting. For all SD experiments, we use the open-source SD v1.4 (Rombach et al., 2022) checkpoint as the pre-trained model. Implementation details and additional results like visualizations of generated images can be found in Appendices B and A.
Baselines. We primarily benchmark against the following baselines commonly used in machine unlearning: (i) Unscrubbed: models trained on data ${\mathcal{D}}$. Unlearning algorithms should scrub information from its parameters. (ii) Finetune (FT) (Golatkar et al., 2020): finetuning models on the remaining data ${\mathcal{D}}_{r}$, ie., catastrophic forgetting. (iii) NegGrad (NG) (Golatkar et al., 2020): gradient ascent on the forgetting data ${\mathcal{D}}_{f}$. (iv) BlindSpot (Tarun et al., 2023b): the state-of-the-art unlearning algorithm for regression. It derives a partially-trained model by training a randomly initialized model with ${\mathcal{D}}_{r}$, then refines the unscrubbed model by mimicking the behavior of this partially-trained model. (v) ESD (Gandikota et al., 2023a): fine-tune the model’s conditional prediction away from the erased concept. (vi) Selective Amnesia (SA) (Heng & Soh, 2023b): adopt EWC from continual learning to preserve model utility when performing forgetting and the method is effective across a wide range of generative models. (vii) SalUn (Fan et al., 2023): the state-of-the-art unlearning algorithm that focuses on salient weights for forgetting across image classification and generation tasks.
Metrics. Several metrics are utilized to evaluate the algorithms: (i) Frechet Inception Distance (FID) (Heusel et al., 2017): the widely-used metric for assessing the quality of generated images. (ii) CLIP score: the similarity between the visual features of the generated image and its corresponding textual embedding. (iii) $P_{\psi}({\mathbf{y}}=c_{f}|{\mathbf{x}}_{f})$ (Heng & Soh, 2023b): the classification rate of a pre-trained classifier $P_{\psi}({\mathbf{y}}|{\mathbf{x}})$, with a ResNet architecture (He et al., 2016) used to classify generated images conditioned on the forgetting classes. A lower classification value indicates superior unlearning performance. (iv) $H(P_{\psi}({\mathbf{y}}|{\mathbf{x}}_{f}))$ (Heng & Soh, 2023b): the average entropy of the classifier’s output distribution given ${\mathbf{x}}_{f}$, which is $H(P_{\psi}({\mathbf{y}}|{\mathbf{x}}_{f}))=-\mathbb{E}[\sum_{i}P_{\psi}({\mathbf{y}}=c_{i}|{\mathbf{x}})\log_{e}P_{\psi}({\mathbf{y}}=c_{i}|{\mathbf{x}})]$. Suppose that all information about $c_{f}$ is erased, the classifier becomes maximally uncertain and the entropy, therefore, would approach $-\sum_{i=1}^{10}\frac{1}{10}\log\frac{1}{10}=2.30$ for CIFAR10.
Table 1: Results on CIFAR10 with DDPM when forgetting the ‘airplane’ class. $H(P_{\psi}({\mathbf{y}}|{\mathbf{x}}_{f}))$ and $P_{\psi}({\mathbf{y}}=c_{f}|{\mathbf{x}}_{f})$ indicate the entropy of the classifier’s distribution and the probability of the forgotten class (ie., the effectiveness of forgetting), respectively. Precision and Recall demonstrate the fidelity and diversity, and FID scores are computed between the generated 45K images and the corresponding ground truth images with the same labels from ${\mathcal{D}}_{r}$ (ie., preserving model utility). The best and the second best are highlighted in blue and orange, respectively.
Unscrubbed FT NG BlindSpot SA SalUn EraseDiff
FID $\downarrow$ 9.63 8.21 76.73 9.12 8.19 8.75 7.61
Precision $\uparrow$ 0.40 0.43 0.08 0.41 0.43 0.43 0.43
Recall $\uparrow$ 0.79 0.77 0.61 0.78 0.75 0.75 0.72
$H(P_{\psi}({\mathbf{y}}|{\mathbf{x}}_{f}))$$\uparrow$ 0.03 0.06 0.45 0.14 1.17 0.04 2.02
$P_{\psi}({\mathbf{y}}=c_{f}|{\mathbf{x}}_{f})$$\downarrow$ 0.97 0.96 0.61 0.90 0.06 0.00 0.22
5.2 Results on DDPM
Following SA, we aim to forget the ‘airplane’ class on CIFAR10. Here, we replace ${\bm{\epsilon}}\in{\mathcal{N}}({\bm{0}},\mathbf{I}_{d})$ with ${\bm{\epsilon}}_{f}\in{\mathcal{U}}({\bm{0}},\mathbf{I}_{d})$. Results are presented in Table 1. Firstly, from Table 1, we can conclude that traditional machine unlearning methods designed for image classification or regression tasks fall short in effectively performing forgetting for DDPM. Finetune and BlindSpot suffer from under-forgetting (ie., the generated image quality is good but the probability of generated images belonging to the forgetting class approaching the value of the unscrubbed model), and NegGrad suffers from over-forgetting (the probability of generated images belonging to the forgetting class is decreased compared to that of the unscrubbed model but the generated image quality drops significantly).
Then, comparing with the unlearning methods SA and SalUn, our algorithm achieves the lowest FID score, indicating that our generated images have the best quality. Note that the FID scores of SA, SalUn, and EraseDiff decrease compared with the generated images from the original models; the quality of the generated images experiences a slight improvement. However, there is a decrease in recall (diversity), which can be attributed to the scrubbed models being fine-tuned over ${\mathcal{D}}_{r}$, suggesting a tendency towards overfitting. Regarding forgetting, SalUn achieves the smallest probability of the generated images classified as the forgetting class; yet, note that the entropy of our algorithm increases significantly, our scrubbed model becomes mostly uncertain about the images conditioned on ${\mathcal{C}}_{f}$, indicating that more information in the generated images conditioned on ${\mathcal{C}}_{f}$ has been erased.
5.3 Results on Stable Diffusion
In this experiment, we apply EraseDiff to forget the ‘tench’ class from Imagenette and erase the ‘nudity’ concept with SD v1.4. For all experiments, we employ SD for sampling with 50 time steps. When forgetting ‘nudity’, we have no access to the training data; instead, we generate $\sim$400 images with the prompts $c_{f}=${‘nudity’, ‘naked’, ‘erotic’, ‘sexual’}.
Forget nudity. 4703 images are generated using I2P prompts, and 1K images are generated using the prompts {‘nudity’, ‘naked’, ‘erotic’, ‘sexual’}. The quantity of nudity content is detected using the NudeNet classifier (Bedapudi, 2019). In Figure 2, the number in the y-axis denotes the number of exposed body parts generated by the SD v1.4 model. Figure 2 presents the percentage change in exposed body parts w.r.t. SD v1.4. In Appendix B, we provide the number of exposed body parts counted in all generated images with different thresholds. Here, our algorithm replaces ${\bm{\epsilon}}_{f}$ with $\epsilon_{{\bm{\theta}}}({\mathbf{x}}_{t}|c_{m})$ where $c_{m}$ is ‘a photo of pokemon’. We can find that, EraseDiff reduces the amount of nudity content compared to SD v1.4, ESD, and SA, particularly on sensitive content like Female/Male Breasts and Female/Male Genitalia. While SalUn excels at forgetting, our algorithm demonstrates a significant improvement in the quality of generated images, as shown in Table 2. Table 2 presented results evaluating the utility of scrubbed models. The FID and CLIP scores are measured over the images generated by the scrubbed models with COCO 30K prompts. While SA achieves the highest CLIP similar score, our algorithm significantly improves the overall quality of the generated images.
Forget class. When performing class-wise forgetting, following Fan et al. (2023), we set the prompt as ‘an image of [$c$]’. For the forgetting class ‘tench’, we choose the ground truth backward distribution to be a class other than ‘tench’. We generate 100 images for each prompt. Note that in Table 3, all unlearning algorithms outperform the original model SD v1.4 in terms of the generated image quality. This could be due to the former methods performing fine-tuning. All methods reduce the probability of classifying the generated images to be the ‘tench’ class to 0, and our proposed algorithm achieves the lowest FID score of 1.29, indicating that EraseDiff can simultaneously perform effectively forgetting and preserving model utility.
Figure 2: Quantity of nudity content detected using the NudeNet classifier from I2P data. Our method effectively erases nudity content from Stable Diffusion (SD), outperforming ESD and SA.
SD v1.4 ESD SA SalUn EraseDiff
FID $\downarrow$ 15.97 15.76 25.58 25.06 17.01
CLIP $\uparrow$ 31.32 30.33 31.03 28.91 30.58
Table 2: Evaluation of generated images by SD when forgetting ‘nudity’. The FID score is measured compared to validation data, while the CLIP similarity score evaluates the alignment between generated images and the corresponding prompts.
SD v1.4 ESD SalUn EraseDiff
FID $\downarrow$ 4.89 1.36 1.49 1.29
$P_{\psi}$$\downarrow$ 0.74 0.00 0.00 0.00
Table 3: Evaluation of generated images by SD when forgetting ‘tench’ from Imagenette. $P_{\psi}$ is short for $P_{\psi}({\mathbf{y}}=c_{f}|{\mathbf{x}}_{f})$ and indicates the probability of the forgotten class (ie., the effectiveness of forgetting).
@@ 5.3 Results on Stable Diffusion
Figure 3: Generated examples with I2P and COCO prompts after forgetting the concept of ’nudity’.
@@
5.4 Computational efficiency
Figure 5: Ablation results.
Memory (MiB) Time (min.)
SA 3352.3 140.00
SalUn 4336.2 28.17
EraseDiff 3360.3 12.70
Figure 6: Cases of potential incomplete erasures.
Finally, we measure the computational complexity of unlearning algorithms. The computational complexity of SA and SalUn involves two distinct stages: the computation of FIM for SA and the computation of salient weights w.r.t. ${\mathcal{D}}_{f}$ for SalUn, and the subsequent forgetting stage for both algorithms. We consider the maximum memory usage across both stages, the metric ‘Time’ is exclusively associated with the duration of the forgetting stage for unlearning algorithms. Table 4 show that EraseDiff outperforms SA and SalUn in terms of efficiency, achieving a speed increase of 11$\sim$ 11$\times$ 11 × than SA and 2$\sim$ 2$\times$ 2 × than SalUn. This is noteworthy, especially considering the necessity for computing FIM in SA for different datasets and models.
5.5 Ablation study
In this experiment, we investigate the influence of the number of iterations $K$ that approximate $\operatorname{min}{\mathcal{L}}_{m}(\bm{\phi};{\mathcal{D}}_{f})$, and the step size $\eta$ that controls the weight of forgetting and preserving model utility. Note that for different hyperparameters in Figure 5, the entropy of the model output remains close to 2.02. This indicates that the scrubbed models become uncertain about the images conditioned on the forgetting class, effectively erasing the information about ${\mathcal{D}}_{f}$. Below, we will further demonstrate the influence on the model utility. In practice, we have $\lambda_{t}=\operatorname{max}\{0,\frac{a_{t}-\nabla_{{\bm{\theta}}}g({\bm{\theta}}_{t})^{\top}\nabla_{{\bm{\theta}}}{\mathcal{L}}_{r}({\bm{\theta}}_{t};{\mathcal{D}}_{r})}{\|\nabla_{{\bm{\theta}}}g({\bm{\theta}}_{t})\|_{2}^{2}}\}=\operatorname{max}\{0,\eta-\frac{\nabla_{{\bm{\theta}}}g({\bm{\theta}}_{t})^{\top}\nabla_{{\bm{\theta}}}{\mathcal{L}}_{r}({\bm{\theta}}_{t};{\mathcal{D}}_{r})}{\|\nabla_{{\bm{\theta}}}g({\bm{\theta}}_{t})\|_{2}^{2}}\}$, we can see that $\eta$ determines the extent to which the update direction for forgetting can deviate from that for preserving model utility. A larger $\eta$ would allow for more deviation in the updating, thus prioritizing forgetting over preserving model utility. In Figure 5, the FID score tends to increase (ie., image quality drop) as the step size $\eta$ increases, indicating that larger $\eta$ leads to greater deviations from the direction that preserves the model utility. Furthermore, the number of iterations $K$ determines how closely the approximation $\bm{\phi}^{K}$ will approach $\operatorname*{arg\,min}_{\bm{\phi}}{\mathcal{L}}_{f}(\bm{\phi};{\mathcal{D}}_{f})$. Hence, a larger number of iterations $K$ leads to more thorough erasure, which is also supported by the results shown in Figure 5, as increasing $K$ correlates with an increase in the FID score.
6 Conclusion and Limitations
In this work, we explored the unlearning problem in diffusion models and proposed an efficient unlearning method EraseDiff. Comprehensive experiments on diffusion models demonstrate the proposed algorithm’s effectiveness in data removal, its efficacy in preserving the model utility, and its efficiency in unlearning. However, our scrubbed model may still preserve some characteristics similar to the forgetting class (e.g., in Figure 6, generated images conditioned on the forgetting class ‘tench’ by our scrubbed model when forgetting the class ‘tench’ from Imagenette, which may preserve some characteristics similar to that close to ‘tench’ visually). Besides, the scrubbed models could be biased for generation, which we do not take into account. Future directions for diffusion unlearning could include assessing fairness post-unlearning, using advanced privacy-preserving training techniques, and advanced MOO solutions. We hope the proposed approach could serve as an inspiration for future research in the field of diffusion unlearning.
Appendix A Details
DDPM.
Results on conditional DDPM follow the setting in SA (Heng & Soh, 2023b). Thanks to the pre-trained DDPM from SA. The batch size is set to be 128, the learning rate is $1\times 10^{-4}$, our model is trained for 200 training steps. 5K images per class are generated for evaluation. For the remaining experiments, four and five feature map resolutions are adopted for CIFAR10 where image resolution is $32\times 32$. All models apply the linear schedule for the diffusion process. We used A5500 and A100 for all experiments.
SD.
We use the open-source SD v1.4 checkpoint as the pre-trained model for all SD experiments. The learning rate is $1\times 10^{-5}$, and our method only fine-tuned the unconditional (non-cross-attention) layers of the latent diffusion model when erasing the concept of nudity. When forgetting nudity, we generate around 400 images with the prompts {‘nudity’, ‘naked’, ‘erotic’, ‘sexual’} and around 400 images with the prompt ‘a person wearing clothes’ to be the training data. We evaluate over 1K generated images for the Imagenette and Nude datasets. 4703 generated images with I2P prompts are evaluated using the open-source NudeNet classifier (Bedapudi, 2019). The repositories we built upon use the CC-BY 4.0 and MIT Licenses.
Appendix B Additional results
Below, we also provide results on SD for EraseDiff when we replace ${\bm{\epsilon}}_{f}$ with $\epsilon_{{\bm{\theta}}}({\mathbf{x}}_{t}|c_{m})$ like Fan et al. (2023); Heng & Soh (2023b), where $c_{m}$ is ‘a person wearing clothes’, denoted as $\text{{EraseDiff}}_{\text{wc}}$. The CLIP score and FID score for $\text{{EraseDiff}}_{\text{wc}}$ are 30.31 and 19.55, respectively.
Table 5: Results on CIFAR10 with conditional DDPM when forgetting the class ‘airplane’. MOO: $\operatorname{min}_{{\bm{\theta}}}{\mathcal{L}}({\bm{\theta}},{\mathcal{D}}_{r})-\lambda{\mathcal{L}}({\bm{\theta}},{\mathcal{D}}_{f})$. ${\mathcal{D}}_{r}^{{}^{\prime}}$: perform unlearning when having no access to the training data. SA outperforms EraseDiff in this scenario.
Unscrubbed MOO EraseDiff EraseDiff (${\mathcal{D}}_{r}^{{}^{\prime}}$) SA (${\mathcal{D}}_{r}^{{}^{\prime}}$)
FID $\downarrow$ 9.63 8.53 7.61 11.74 9.63
Precision $\uparrow$ 0.40 0.41 0.43 0.38 0.40
Recall $\uparrow$ 0.79 0.76 0.72 0.76 0.77
$H(P_{\psi}({\mathbf{y}}|{\mathbf{x}}_{f}))$$\uparrow$ 0.03 2.02 2.02 1.59 0.93
$P_{\psi}({\mathbf{y}}=c_{f}|{\mathbf{x}}_{f})$$\downarrow$ 0.97 0.21 0.22 0.36 0.03
Figure 8: Quantity of nudity content detected using the NudeNet classifier with a threshold of 0.6.
Figure 9: Quantity of nudity content detected using the NudeNet classifier with a threshold of 0.4.
Figure 10: Quantity of nudity content detected using the NudeNet classifier with a threshold of 0.2.
Figure 11: Generated examples with I2P prompts when forgetting the concept of ’nudity’.
Figure 12: Generated examples with I2P prompts when forgetting the concept of ’nudity’.
Figure 13: Generated examples with I2P prompts when forgetting the concept of ’nudity’.
Figure 15: Visualization of generated examples with prompts {‘nudity’, ‘naked’, ‘erotic’, ‘sexual’} when forgetting the concept of ’nudity’.
Figure 16: Visualization of generated images with COCO 30K prompts by the scrubbed SD models when forgetting the concept of ’nudity’.
Figure 17: Visualization of generated images with COCO 30K prompts by the scrubbed SD models when forgetting the concept of ’nudity’.

### v4

@@ 7 Proofs
The objective aims to minimize $\mathcal{L}_{r}({\bm{\theta}};{\mathcal{D}}_{r})+\lambda g({\bm{\theta}};\bm{\phi}^{\ast}({\bm{\theta}}))$, let ${\bm{\theta}}^{\ast}$ be the optimal solution to this objective. Note that $g({\bm{\theta}},\bm{\phi}^{\ast}({\bm{\theta}}))={\mathcal{L}}_{f}({\bm{\theta}};{\mathcal{D}}_{f})-\operatorname{min}{\mathcal{L}}_{f}(\bm{\phi}^{\ast}({\bm{\theta}});{\mathcal{D}}_{f})\geq 0$ as $\bm{\phi}$ starts from ${\bm{\theta}}$ and is update to decreas ${\mathcal{L}}_{f}(\bm{\phi};{\mathcal{D}}_{f})$. This will decrease to 00 for minimizing the above objective. Therefore, at the optimal solution ${\bm{\theta}}^{\ast}$, we have $g({\bm{\theta}}^{\ast},\bm{\phi}^{\ast}({\bm{\theta}}^{\ast}))=0$. This further implies that ${\mathcal{L}}_{f}({\bm{\theta}}^{*};{\mathcal{D}}_{f})=\operatorname{min}{\mathcal{L}}_{f}(\bm{\phi}^{\ast}({\bm{\theta}}^{*});{\mathcal{D}}_{f})$, meaning that ${\bm{\theta}}^{*}$ is the current optimal solution of ${\mathcal{L}}_{f}({\bm{\theta}};D_{f})$ because we cannot update further the optimal solution. Moreover, we have ${\bm{\theta}}^{*}$ as the local minima of $\mathcal{L}_{r}({\bm{\theta}};{\mathcal{D}}_{r})$ in sufficiently small vicinity considered, because in the small vicinity around ${\bm{\theta}}^{*}$, $g\left({\bm{\theta}},\bm{\phi}^{*}({\bm{\theta}}^{*})\right)=0$ provides no further improvements for the above sum, any increase in the above objective in the vicinity of ${\bm{\theta}}^{*}$ would primarily be due to an increase in $\mathcal{L}_{r}({\bm{\theta}};{\mathcal{D}}_{r})$.
8 Reproducibility Statement and Details
In this section, we provide detailed instructions on the reproduction of our results, we also share our source code at the repository https://github.com/JingWu321/EraseDiff.
DDPM.
Results on conditional DDPM follow the setting in SA [28]. Thanks to the pre-trained DDPM from SA. The batch size is set to be 128, the learning rate is $1\times 10^{-4}$, our model is trained for around 300 training steps. 5K images per class are generated for evaluation. For the remaining experiments, four and five feature map resolutions are adopted for CIFAR10 where image resolution is $32\times 32$. All models apply the linear schedule for the diffusion process. We used A5500 and A100 for all experiments.
SD.
We use the open-source SD v1.4 checkpoint as the pre-trained model for all SD experiments. The learning rate is $1\times 10^{-5}$, and our method only fine-tuned the unconditional (non-cross-attention) layers of the latent diffusion model when erasing the concept of nudity. When forgetting nudity, we generate around 400 images with the prompts {‘nudity’, ‘naked’, ‘erotic’, ‘sexual’} and around 400 images with the prompt ‘a person wearing clothes’ to be the training data. We evaluate over 1K generated images for the Imagenette and Nude datasets. 4703 generated images with I2P prompts are evaluated using the open-source NudeNet classifier [2]. The repositories we built upon use the CC-BY 4.0 and MIT Licenses.
9 Additional results
Below, we also provide results on SD for EraseDiff when we replace ${\bm{\epsilon}}_{f}$ with $\epsilon_{{\bm{\theta}}}({\mathbf{x}}_{t}|c_{m})$ like Fan et al. [16], Heng and Soh [28], where $c_{m}$ is ‘a person wearing clothes’, denoted as $\text{{EraseDiff}}_{\text{wc}}$. The CLIP score and FID score for $\text{{EraseDiff}}_{\text{wc}}$ are 30.31 and 19.55, respectively.
Table 5: Performance of class-wise forgetting on Imagenette using SD. UA: the accuracy of the generated images that do not belong to the forgetting class (i.e., the effectiveness of forgetting). The FID score is measured compared to validation data for the remaining classes.
Forget. Class SalUn EraseDiff
FID $\downarrow$ UA (%$\uparrow$ FID $\downarrow$ UA (%$\uparrow$
Tench 1.49 100 1.29 100
English Springer 1.50 100 1.38 100
Cassette Player 1.11 100 0.85 100
Chain Saw 1.64 100 1.17 99.9
Church 0.76 100 0.83 100
French Horn 0.67 100 1.09 100
Garbage Truck 1.54 100 0.96 100
Gas Pump 1.59 100 1.25 100
Golf Ball 1.29 98.8 1.50 99.5
Parachute 1.35 100 0.78 99.7
Average 1.29 99.88 1.11 99.91
@@ 9 Additional results
Figure 7: Generated examples with I2P prompts when forgetting the concept of ‘nudity’.
Figure 8: Generated examples with I2P prompts when forgetting the concept of ‘nudity’.
Figure 9: Generated examples with I2P prompts when forgetting the concept of ‘nudity’.
Figure 11: Visualization of generated examples with prompts {‘nudity’, ‘naked’, ‘erotic’, ‘sexual’} when forgetting the concept of ‘nudity’.
Figure 12: Visualization of generated images with COCO 30K prompts by the scrubbed SD models when forgetting the concept of ‘nudity’.
Figure 13: Visualization of generated images with COCO 30K prompts by the scrubbed SD models when forgetting the concept of ‘nudity’.